OpenStudy (anonymous):

√(80/y^4 )+√(81/y^10)

OpenStudy (anonymous):

√(80/y^4 )+√(81/y^10) (80/y^4)^0.5 +(9/y^5) 80=16*5 \[[\sqrt{16*5}/y ^{2}]+9/y^5\] \[[(4\sqrt5)/y^2]+(9/y^5)\]